Help with these problems tutors please thanks!

A poll was taken of 13,604 working adults aged 40-70 to determine their level of education. The participants were classified by sex and by level of education. The results are shown below.
Education Level Male Female Total
High School or Less 2700 3030 5730
Bachelor's Degree 2800 3993 6793
Master's Degree 473 502 975
Ph.D. 60 46 106
Total 6033 7571 13,604

A person is selected at random. Compute the following probabilities.
(a) What is the probability that the selected person does not have a Ph.D.?

(b) What is the probability that the selected person does not have a Master's degree?

(c) What is the probability that the selected person is female or has a Master's degree?

(d) What is the probability that the selected person is male or has a Ph.D.?

Answer by ewatrrr(24785)


Hi
Education Level Male Female Total
High School or Less 2700 3030 5730
Bachelor's Degree 2800 3993 6793
Master's Degree 473 502 975
Ph.D. 60 46 106
Total 6033 7571 13,604
A person is selected at random. Compute the following probabilities.
You did not say Decimal 0r Fraction????
(a) What is the probability that the selected person does not have a Ph.D.?
P = 1 - 106/13604 = 13498/13604 = 6749/6802
(b) What is the probability that the selected person does not have a Master's degree?
P = 1 - 975/13604 = 12629/13604
(c) What is the probability that the selected person is female or has a Master's degree?
P = (7571 + 975 - 502) / 13604 = = 8044/13604= 2011/3401
(d) What is the probability that the selected person is male or has a Ph.D.?
P = (6033 + 106 - 60)/13604 = 6079/13604